The ratio of the measure of a supplement of an angle to the measure of a complement is 4:1. Find the measure of the angle, the complement and the supplement.
Answer by mananth(16949)

let angle be x
..
supplement will be 180-x
complement will be 90-x
..
180-x/(90-x)=4/1
180-x=4(90-x)
180-x=360-4x
-x+4x=360-180
3x=180
/3
x=60 deg. the measure of the angle.
...
m.ananth@hotmail.ca